(1 point) How many rounds of golf do those physicians who play golf play per year? A survey of 12 physicians revealed the folThanks

Will upvote!
Answer #1

Mean = 24.08333


Confidence Interval :-
X̅ ± Z( α /2) σ / √ ( n )
Z(α/2) = Z (0.03 /2) = 2.170
24.0833 ± Z (0.03/2 ) * 6/√(12)
Lower Limit = 24.0833 - Z(0.03/2) 6/√(12)
Lower Limit = 20.3248
Upper Limit = 24.0833 + Z(0.03/2) 6/√(12)
Upper Limit = 27.8419
97% Confidence interval is ( 20.3248 , 27.8419 )